OpenStudy (anonymous):

Choose the correct answer.________Yves bought 420 tropical fish for a museum display. He bought 6 times as many parrotfish as angelfish. How many of each type of fish did he buy? Which is a system of equations to model the problem if x represents the number of angelfish Yves bought and y represents the number of parrotfish he bought? A. x + y = 6 y = 420x __________________ B. x + 6y = 420 y = 6x ______________ C. x – y = 420 y = 6x _________________ D. x + y = 420 y = 6x

OpenStudy (anonymous):

@Mehek14 @LieutenantGeneral @fomf @girlmeetsworld

OpenStudy (anonymous):

@DavidUsa

OpenStudy (anonymous):

help @AskMeLotss

OpenStudy (anonymous):

I dont know this :( sorry

OpenStudy (anonymous):

@Michele_Laino @Destinymasha @Nnesha

OpenStudy (michele_laino):

If I call with \(p\) the number of parrotfish and with \(a\) the number of angelfish, then we can note that: the statement "He bought 6 times as many parrotfish as angelfish" means that \(p=6a\) and the statement: "Yves bought 420 tropical fish" means \(p+a=420\)

OpenStudy (michele_laino):

@Cadance.r

OpenStudy (anonymous):

Ok I am working it out @Michele_Laino

OpenStudy (anonymous):

@Michele_Laino would it be B x+6y=240 y=6x

OpenStudy (michele_laino):

no, I'm sorry, the second equation is right, whereas the first one is wrong. please keep in mind that \(420\) is the total number of fish, so we can write \(x+y=420\)

OpenStudy (anonymous):

Wait a minute let me recalculate

OpenStudy (anonymous):

@Michele_Laino Is it C x – y = 420 y = 6x

OpenStudy (michele_laino):

I think that the first equation is \(x+y=420\) whereas, the second equation is \(y=6x\) so, what is the right option?

OpenStudy (michele_laino):

total number of fish is \(x+y\) not \(x-y\)

OpenStudy (anonymous):

SO IT IS D @Michele_Laino

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Yya thank you thank thank you so much @Michele_Laino I got a few more Questions to do.Can you help me on them?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

I can help!

OpenStudy (lieutenantgeneral):

IM GOING WITH D. THIS IS EASY FIRST DIVIDE see what ya get and it says x = blah blah and y = blah soo it's adding too x + y = 420 DOI